# Direct probe of magnetic field effects on phonons by ultrasound propagation in a quasi-two-dimensional honeycomb magnet Na2Co2TeO6

### v1

@@ IV Summary
To summarize, using the ultrasound technique we studied the temperature and field induced transitions in NCTO. We observed strongly asymmetric and hysteretic features in the acoustic properties of NCTO in the magnetic field applied perpendicular to the honeycomb plane in opposite directions. Quantitative analyses of the hysteresis area unveiled a sharp maximum reached at around 12 K, with a strong suppression when the temperature is further lowered. This observation suggests the scenario with triple-Q ground state in NCTO rather than a zigzag ground state proposed from an XXZ model. Our ultrasound results are in line with the thermal transport data known from the literature Yang2022 ; Hong2024 . This fact strongly suggests that the intricate thermal transport features of NCTO at low temperatures have no direct contribution from novel excitations beside phonons. Our results highlight that phonon effects must be considered very carefully when evaluating the experimental data of frustrated magnets.
@@ V Acknowledgments
We thank Jérémy Sourd and Andreas Hauspurg for technical support. This work has been supported by the Deutsche Forschungsgemeinschaft (DFG) through Project No. 247310070 (SFB 1143), Project No. 390858490 (Würzburg-Dresden Cluster of Excellence ct.qmat, EXC 2147), and Project No. 411750675 (Emmy Noether program, JA2306/4-1). This project has received funding from the European Research Council (ERC) under the European Unions-Horizon 2020 research and innovation programme (grant agreement No. 647276-MARS-ERC-2014-CoG). We acknowledge support of the HLD at HZDR, member of the European Magnetic Field Laboratory (EMFL). XCH acknowledge the start-up funding of Chongqing University.

### v2

@@ IV Summary
To summarize, using the ultrasound technique we studied the temperature and field induced transitions in NCTO. We observed strongly asymmetric and hysteretic features in the acoustic properties of NCTO in the magnetic field applied perpendicular to the honeycomb plane in opposite directions. Quantitative analyses of the hysteresis area unveiled a sharp maximum reached at around 12 K, with a strong suppression when the temperature is further lowered. This observation suggests the scenario with triple-Q ground state in NCTO rather than a zigzag ground state proposed from an XXZ model. A comparative analysis of our ultrasound measurements with previously reported thermal transport data reveals the pivotal role of phonons in governing the complex thermal transport properties of NCTO at low temperatures Yang2022 ; Hong2024 ; Li2023 . These findings underscore the necessity of carefully accounting for phonon-related effects when interpreting experimental results in frustrated magnetic systems.
@@ V Acknowledgments
We thank Jérémy Sourd and Andreas Hauspurg for technical support. We thank Robin Neumann for valuable feedback on the manuscript and for sharing their experimental data with us. This work has been supported by the Deutsche Forschungsgemeinschaft (DFG) through Project No. 247310070 (SFB 1143), Project No. 390858490 (Würzburg-Dresden Cluster of Excellence ct.qmat, EXC 2147), and Project No. 411750675 (Emmy Noether program, JA2306/4-1). This project has received funding from the European Research Council (ERC) under the European Unions-Horizon 2020 research and innovation programme (grant agreement No. 647276-MARS-ERC-2014-CoG). We acknowledge support of the HLD at HZDR, member of the European Magnetic Field Laboratory (EMFL). XCH acknowledge the start-up funding of Chongqing University.
VI DATA AVAILABILITY
The data that support the findings of this article are openly available Data .